OpenStudy (anonymous):

Find dy/dx by implicit differentiation if \[(x^2+y^2)^2-(\pi)x^2y=0\]

OpenStudy (adi3):

@ganeshie8

OpenStudy (perl):

you can use the chain rule

OpenStudy (perl):

$$ \large 2(x^2 + y^2) (2x + 2y ~ y ~' )- \pi(2x y + x^2~ y ' ) = 0 $$

OpenStudy (anonymous):

okay

OpenStudy (perl):

you can also write it as $$ \large 2(x^2 + y^2) (2x + 2y ~ \dfrac{dy}{dx} )- \pi(2x y + x^2\frac{dy}{dx} ) = 0 $$

OpenStudy (anonymous):

what do i do after? @perl

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@rational

ganeshie8 (ganeshie8):

solve \(y'\)

OpenStudy (anonymous):

how do we solve y'?

ganeshie8 (ganeshie8):

\[\large 2(x^2 + y^2) (2x + 2y ~ \color{red}{y '} )- \pi(2x y + x^2~ \color{red}{y '} ) = 0\]

ganeshie8 (ganeshie8):

you're done with calculus part... isolating \(\color{red}{y'}\) is just algebra

OpenStudy (anonymous):

how not sure how...

OpenStudy (anonymous):

do i expand it?

ganeshie8 (ganeshie8):

yes

ganeshie8 (ganeshie8):

\[2(x^2 + y^2) (2x + 2y ~ \color{red}{y '} )- \pi(2x y + x^2~ \color{red}{y '} ) = 0\] \[2(x^2 + y^2) (2x) + 2(x^2 + y^2) (2y ~ \color{red}{y '} )- \pi(2x y) -\pi( x^2~ \color{red}{y '} ) = 0\]

OpenStudy (anonymous):

how did you that?

ganeshie8 (ganeshie8):

distributed

ganeshie8 (ganeshie8):

\[(a+b)(c+d) = (a+b)c + (a+b)d\]

OpenStudy (anonymous):

how did you get c and d

ganeshie8 (ganeshie8):

that was just an example

OpenStudy (anonymous):

(2x) and (2y y')?

ganeshie8 (ganeshie8):

You have this equation to start with : \[(x^2+y^2)^2-(\pi)x^2y=0\] implicitly differentiate with respect to x, you get : \[2(x^2 + y^2) (2x + 2y ~ \color{red}{y '} )- \pi(2x y + x^2~ \color{red}{y '} ) = 0\] you get this far ? :)

OpenStudy (anonymous):

yes

ganeshie8 (ganeshie8):

good, next expand everything. what do you get ?

OpenStudy (anonymous):

\[2(2x^3+2x^2y y'+2xy'+y^2y')-(\pi)2xy+(\pi)x^2y'=0\] i got this, it looks weird.

OpenStudy (anonymous):

oh i see, what you were doing. You first distribute the entire first term to the another.

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

yes

ganeshie8 (ganeshie8):

\[2(x^2 + y^2) (2x + 2y ~ \color{red}{y '} )- \pi(2x y + x^2~ \color{red}{y '} ) = 0\] \[2(x^2 + y^2) (2x) + 2(x^2 + y^2) (2y ~ \color{red}{y '} )- \pi(2x y) -\pi( x^2~ \color{red}{y '} ) = 0\] collect \(\color{red}{y'}\) terms \[2(x^2 + y^2) (2x) - \pi(2x y) + \color{red}{y '} [2(x^2 + y^2) (2y ~ )-\pi( x^2)] = 0\] \[ \color{red}{y '} [2(x^2 + y^2) (2y ~ )-\pi( x^2)] = \pi(2x y) -2(x^2 + y^2) (2x)\] \[ \color{red}{y '} = \dfrac{ \pi(2x y) -2(x^2 + y^2) (2x)}{ 2(x^2 + y^2) (2y ~ )-\pi( x^2)}\]

ganeshie8 (ganeshie8):

simplify if you want to

OpenStudy (anonymous):

So what you did to collect y' terms, you common factor? @ganeshie8

OpenStudy (anonymous):

i'm confused. How to simplify? The answer from the text says: \[\frac{ 2x((\pi)y-2x^2-2y^2) }{ 4x^2y+4y^3-(\pi)x^2 }\]

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\[\color{red}{y '} = \dfrac{ \pi(2x y) -2(x^2 + y^2) (2x)}{ 2(x^2 + y^2) (2y ~ )-\pi( x^2)}\] expand terms in the denominator

OpenStudy (anonymous):

i got the denominator, i can't get the numerator

ganeshie8 (ganeshie8):

\[\color{red}{y '} = \dfrac{ \pi(\color{purple}{2x} y) -2(x^2 + y^2) (\color{purple}{2x})}{ 2(x^2 + y^2) (2y ~ )-\pi( x^2)}\] factor out \(\color{purple}{2x}\) from numerator

OpenStudy (anonymous):

you can do that?

ganeshie8 (ganeshie8):

why not

OpenStudy (anonymous):

it looks strange to me because (2xy) are together?

ganeshie8 (ganeshie8):

\[\large a\heartsuit \spadesuit + b\heartsuit = \heartsuit(a\spadesuit + b)\]

OpenStudy (anonymous):

oh, i see your point :)

OpenStudy (anonymous):

thank you too much for helping me and answering my silly questions!

ganeshie8 (ganeshie8):

yw:)

ganeshie8 (ganeshie8):

there are no silly questions, only silly answers :P